### v1

# Dexterous Manipulation through Imitation Learning: A Survey

Abstract
Dexterous manipulation, which refers to the ability of a robotic hand or multi-fingered end-effector to skillfully control, reorient, and manipulate objects through precise, coordinated finger movements and adaptive force modulation, enables complex interactions similar to human hand dexterity. With recent advances in robotics and machine learning, there is a growing demand for these systems to operate in complex and unstructured environments. Traditional model-based approaches struggle to generalize across tasks and object variations due to the high-dimensionality and complex contact dynamics of dexterous manipulation. Although model-free methods such as reinforcement learning (RL) show promise, they require extensive training, large-scale interaction data, and carefully designed rewards for stability and effectiveness. Imitation learning (IL) offers an alternative by allowing robots to acquire dexterous manipulation skills directly from expert demonstrations, capturing fine-grained coordination and contact dynamics while bypassing the need for explicit modeling and large-scale trial-and-error. This survey provides an overview of dexterous manipulation methods based on imitation learning (IL), details recent advances, and addresses key challenges in the field. Additionally, it explores potential research directions to enhance IL-driven dexterous manipulation. Our goal is to offer researchers and practitioners a comprehensive introduction to this rapidly evolving domain.
Index Terms:
Dexterous Manipulation, Imitation Learning, End Effector, Teleoperation
I Introduction
Over the past few decades, robotics has attracted intensive research interests, with dexterous manipulation emerging as a particularly popular focus. Dexterous manipulation aims to perform complex, precise and flexible tasks (such as grasping an object, opening a drawer, and rotating a pencil) in various scenes with human-level dexterity using robotic hands or other end-effectors. This high-precision manipulation capability supports a board spectrum of applications, including industrial manufacturing [1, 2, 3, 4], space or underwater exploration, [5, 6, 7, 8] and medical care [9, 10, 11, 12]. Recently, the rapid development of imitation learning [13, 14], which seeks to acquire knowledge by observing and mimicking behaviors of humans or other agents, has led to notable advancements in computer graphics and robotics. As an intuitive approach to equip robots with human prior knowledge, especially in the ability to interact with objects and understand the scenes, IL has shown exceptional performance in enabling robots to perform tasks with human-like dexterity.
Figure 1: Examples of dexterous manipulation in the real world. Row 1: Our captured image, [15], [16], Row 2: [17], [18], Row 3: [19].
Figure 2: Overview of dexterous manipulation based on imitation learning in this survey.
Research on the dexterous manipulation has received significant attention even before reinforcement learning (RL) was adopted to optimize their behavior strategies through iterative interactions with the environment and reward-based feedback mechanisms. Traditional approaches encourage robots acquire dexterous manipulation skills by modeling domain dynamics and applying optimal control methods. These approaches are theoretically sound but rely heavily on the fidelity of the world model.
However, when it comes to dexterous operations, such as assembling precision components or performing complex surgical procedures, a high degree of flexibility and multi-degree-of-freedom movement capabilities are required to execute complex, human-like tasks. The successful execution of dexterous manipulation hinges upon intricate and highly precise mechanical design, such as multi-fingered robotic hand [20] or anthropomorphic arms [21], as well as sophisticated control algorithms required to handle high-dimensional spaces [21] and multi-contact dynamics [22].
Recently, the exploration of employing IL to the field of robotics has garnered significant attention from researchers. Without the need for crafting complex world models and carefully designed reward functions, IL enables robots to learn tasks by observing and imitating expert demonstrations. This approach is intuitive, as the goal is for robots to substitute human labor by performing tasks like human experts. Specifically, the initial step involves collecting a dataset of expert demonstrations, which contains trajectories of manipulation tasks conducted by humans or well-trained agents. Robots use such trajectories as a reference to learn task behaviors. To ensure consistency, it is preferable to use identical robots during both data collection and execution phases. However, this implementation is not conducive to facilitating data sharing among heterogeneous robotic systems. One solution is to map the trajectory of original manipulator to the target robots, a process known as retargeting. Nonetheless, the process of humans operating robots for data collection remains a time-consuming and labor-intensive task in the context of constructing large-scale datasets. To address this issue, researchers[23, 24, 17] have adopted pose estimation techniques from computer vision to develop mappings from human hands to robotic hands, effectively lowering the barriers of collecting demonstration data. Additionally, dataset augmentation enhances the ability to generalize to new objects and scenes, contributing to the expansion of the dataset.
IL mimics expert behaviors analogous to supervised learning (SL), and often integrates with reinforcement learning (RL) to address complex decision-making tasks. Both IL and SL share similarities in learning from demonstrations or ground truth data. However, their objectives differ: SL aims to produce outputs identical to the ground truth in static scenarios, while IL focuses on task completion in dynamic environments, such as changes in target object position and environmental disturbances in manipulation tasks. Such tasks usually involves sequential decision-making, where errors can accumulate over steps, leading to compounding errors and overall task failure. IL emphasizes task completion by adjusting and compensating for initial errors in subsequent decisions, thereby reducing their overall impact. In dexterous manipulation tasks, RL and Il are often combined. This combination addresses the inefficiencies caused by the agent’s large and complex action space has a high degree of freedom and makes pure RL exploration less effective. IL leverages expert demonstrations to offer an straight guidance, thereby reducing exploration time and increasing efficiency. Additionally, reward functions for manipulation tasks are often challenging to design, different tasks typically necessitate distinct reward functions. However, IL benefits from a relatively universal reward function to fit demonstration trajectories, necessitating only the provision of varied demonstration data, which ultimately improves learning efficiency and task success rates.
IL is particularly advantageous for dexterous manipulation tasks. This is because objects involved in dexterous manipulation are typically designed for human use, and thus robots subject to these tasks are likely to have structures similar to humans or parts of the human body, such as humanoid robots, dual-arm manipulators, or dexterous hands. They usually require precise control, coordination, and adaptability, attributes that are challenging to achieve through traditional methods. Specifically, this learning paradigm includes various branches such as behavior cloning [25], [26], hybrid approaches (the combination of RL and IL) [27], [28], hierarchical IL [29], [30] and others, each contributing unique advantages to the learning process.
Since the intersection of IL and dexterous manipulation represents a frontier in robotics research. In the past decade, several works including DAPG [27], which combines deep reinforcement learning(DRL) with human demonstrations to solve high-dimensional dexterous manipulation tasks; Implicit Behavioral Cloning [26] which focuses on improving robot policy learning from a mathematical perspective; Hiveformer [31], which explores creating multimodal interactive agents; Diffusion Policy [32] leverages recent advancement in generative models to achieve better performance in manipulation tasks, have been proposed and significantly expanded the boundaries of what is achievable in robotic dexterous manipulation. However, despite the recent considerable progress in this field, numerous challenges remain. Data collection for IL is labor-intensive and time-consuming [33], [16]. The acquisition of generalization ability from learned behaviors to new tasks and varying environments is also non-trivial [34]. Additionally, real-time control and sim-to-real transfer, where robots trained in simulation must perform effectively in the real world, both hinder the application [35]. Addressing these challenges requires a concerted effort in developing more efficient data collection methods, improving learning algorithms, and enhancing the physical capabilities of robotic systems.
The main purpose of this survey is to provide an overview about IL-based dexterous manipulation approaches. The rest of this article is organized as follows: In Section II, an introduction to both dexterous manipulation and IL is presented in detail. Subsequently, we discuss the state-of-the-art IL-based dexterous manipulation techniques and highlight notable achievements in this field in Section III. Section IV discusses end-effectors for dexterous manipulation. Moreover, we discuss teleoperation systems and datasets in Section V VI. We summarize existing challenges and propose future directions for research in this rapidly evolving field in Section VII and VIII respectively. Finally, conclusions are made in Section IX. By synthesizing the existing body of knowledge, this survey aspires to serve as a valuable resource for researchers and practitioners seeking to advance the capabilities of robotic systems through the synergy of IL and dexterous manipulation.
II Overview of Imitation Learning based Dexterous Manipulation
II-A Dexterous Manipulation
In the field of robotics, dexterous manipulation [36, 37, 38] refers to the capability of robotic systems to execute intricate and precise tasks. These tasks often employ grippers or dexterous hands to grasp, maneuver, and manipulate objects [39]. Characterized by high degrees of freedom and fine motor skills, dexterous manipulation extends beyond simple pick-and-place operations to include activities such as tool use, object reorientation, and complex assembly tasks. Achieving such manipulation commonly involves using sophisticated end-effectors designed to emulate the versatility and finesse of human hands, such as multi-fingered hands or anthropomorphic robotic arms.
Dexterous manipulation poses several significant challenges, including precise control, high-dimensional motion planning, and real-time adaptability to dynamic environments [40]. The intricacies of these tasks demand not only robust mechanical design but also advanced control algorithms capable of handling the complexities of multi-contact interactions and the variability inherent in real-world scenarios.
Traditional model-based methods [41, 42] have become inadequate for robots performing complex tasks due to the increasing complexity of manipulation tasks. Consequently, extensive research has been dedicated to learning-based approaches, with reinforcement learning (RL) emerging as an effective method. Various works have exploited RL for robots to learn dexterous policies [43, 44, 45, 46, 47]. However, pure RL has several inherent drawbacks. RL algorithms often struggle with exploring high-dimensional action spaces efficiently in dexterous manipulation tasks [48]. Additionally, designing reasonable reward functions is challenging; flawed reward functions can affect exploration and learning speed, leading to inferior performance.
Recently, advancements in imitation learning (IL) have opened new avenues for addressing these challenges.
II-B Imitation Learning
The main purpose of IL is to enable agents to learn and perform behaviors by imitating expert demonstrations [35]. In contrast, pure RL requires carefully designed reward functions and is particularly effective in scenarios where the desired behavior is difficult to describe in algorithms but can be easily demonstrated. IL employs these expert demonstrations to guide the learning process of agents by establishing the correlation between observed states and corresponding actions. Through IL, agents can transcend merely replicating basic and predefined behaviors within controlled and constrained environments, enabling them to autonomously execute optimal actions in complex, unstructured environments [49]. Consequently, IL significantly alleviates the burden on experts, facilitating efficient skill transfer.
Methodologies of IL can be broadly categorized into several sub-classes, including behavior cloning [50], inverse reinforcement learning(IRL) [51], and generative adversarial imitation learning(GAIL) [52]. Behavior cloning directly maps observed actions to the agent’s actions through SL techniques. IRL, on the other hand, aims to deduce the underlying reward structure that motivates the demonstrator’s behavior, allowing the agent to optimize its actions accordingly. GAIL employs adversarial training techniques to improve the imitation policy by distinguishing between expert and agent actions, thus refining the agent’s ability to replicate the desired behavior accurately.
TABLE I: Comparison of different imitation learning approaches.
Approach Key Characteristics Pros Cons
Behavioral Cloning [53, 54, 20, 55, 14, 56, 57, 29, 53, 26, 58, 59, 60, 32, 61, 62] Supervised learning paradigm. Direct mapping from states to actions. No reward signals or exploration. Simple and easy to implement. Data-efficient with large
demonstrations.
Prone to distribution shift. Poor generalization to
unseen states.
Inverse Reinforcement
Learning
[63, 64, 65, 66, 67, 68, 45, 69, 70, 71, 72]
Inferring expert’s reward function. Deriving policy by maximizing the
inferred reward.
Generalization to new situations. Suitable for tasks with unknown
rewards.
Computationally intensive. Non-unique reward solutions.
Generative Adversarial
Imitation Learning
[73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87]
Adversarial training between
generator and discriminator.
No explicit reward function.
Good sample efficiency. Improved robustness. Training instability. Mode collapse and sensitivity
to hyperparameters.
Hierarchical Imitation
Learning
[88, 89, 90, 91, 92, 93, 94]
Two-level hierarchical policy. Decomposing tasks into sub-tasks
and primitives.
Scalable to complex tasks. Modular and reusable
sub-policies.
Requiring hierarchy design. Requiring training coordination.
Continual Imitation
Learning
[95, 96, 92, 97, 98, 99]
Continual skill acquisition. Adapting previously learned
behaviors.
Flexible to evolving tasks. Reducing forgetting of old skills. Risk of catastrophic forgetting. Requiring ongoing expert input.
III Imitation Learning based Dexterous Manipulation Approaches
We categorize IL-based dexterous manipulation approaches into four categories: (1) Behavioral Cloning, (2) Inverse Reinforcement Learning, (3) Generative Adversarial Imitation Learning, and other extended frameworks, including (4) Hierarchical Imitation Learning and (5) Continual Imitation Learning. In the following subsections, we provide an overview of each category, followed by a detailed description and a summary of key research progress. Tab. I presents a comparison between different imitation learning approaches.
III-A Behavioral Cloning
III-A1 Description
Behavioral Cloning (BC) refers to replicating expert behavior by learning directly from demonstrated state-action pairs. Specifically, BC is characterized by (1) a supervised learning paradigm and (2) a direct mapping from states to actions without relying on reward signals or exploration, as is typical in RL.
To formally define BC, we consider a set of $n$ demonstrations $\mathcal{D}=\{\tau_{1},\dots,\tau_{n}\}$, where each demonstration $\tau_{i}$ is a sequence of state-action pairs of length $N_{i}$. Specifically, $\tau_{i}=\{(s_{1},a_{1}),\dots,(s_{N_{i}},a_{N_{i}})\}$, with states $s\in\mathcal{S}$ and actions $a\in\mathcal{A}$. $\mathcal{S}$ and $\mathcal{A}$ denote the state and action spaces, respectively. The objective of BC is to learn a policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$ that imitates the expert behavior by minimizing the negative log-likelihood of the demonstrated actions. Formally, the objective function is:
$\mathcal{L}(\pi)=-\mathbb{E}_{(s,a)\sim p_{\mathcal{D}}}\left[\log\pi(a\mid s)\right]$ (1)
III-A2 Research Progress
BC has achieved significant progress in dexterous manipulation [53, 54, 100, 101, 20] and has demonstrated effective performance in relatively simple tasks, such as pushing [55] and grasping [14]. However, its applicability in dynamic environments and long-horizon tasks remains an active area of research.
The training data for BC models are usually collected from expert demonstrations tailored to specific tasks. Consequently, when the agent encounters states that are unseen during training, it may produce actions that deviate from the expert’s behavior, leading to task failure. In sequential decision-making processes, even small deviations from expert actions at each step can accumulate over time, resulting in what is known as the “compounding error” problem. This issue is particularly pronounced in dexterous manipulation tasks[56, 57], due to the high dimensionality of the action space and the strong dependency between task success and the consistency of the predicted action trajectory. To mitigate compounding errors in dexterous manipulation, Mandlekar et al. [29] propose a hierarchical framework that segments demonstration trajectories at intersection points across different tasks and recombines them to synthesize trajectories for unseen tasks. Similarly, Zhao et al. [53] address the problem by considering the compatibility with high-dimensional visual observations. Instead of predicting actions step by step, they propose to predict entire action sequences, thereby reducing the effective decision horizon and alleviating compounding errors.
Another challenge in BC is its limited ability to model multi-modal data, which is prevalent in human demonstrations collected from real-world environments. To overcome this limitation, several approaches have been proposed to model multi-modal action distributions. Florence et al.[26] formulate BC as a conditional energy-based modeling problem for capturing multi-modal data distribution, albeit at the cost of increased computational overhead. Similarly, Shafiullah et al.[58] propose to model the action distribution as a mixture of Gaussians. Their method leverages the Transformer architecture to efficiently utilize the history of previous observations and enables multi-modal action prediction through token-based outputs. Another promising direction is leveraging generative models to capture the inherent diversity of expert behaviors. Mandlekar et al.[59] propose using generative models for trajectory prediction, enabling selective imitation, though this approach relies on carefully curated, task-specific datasets.
More recently, diffusion models have shown great potential in enhancing the robustness and generalization of BC methods. Chen et al. [60] propose a diffusion-augmented BC framework that models both conditional and joint probability distributions over expert demonstrations. Building on this idea, Chi et al. [32] employ diffusion models as decision models to directly generate sequential actions conditioned on visual input and the robot’s current state. Additionally, the 3D Diffusion Policy [61] leverages 3D input representations to better capture scene spatial configurations. Similarly, the 3D Diffuser Actor [62] utilizes full 3D scene representations by integrating RGB and depth information, along with language instructions, robot proprioception, and noise trajectories, through a 3D Relative Transformer framework.
III-A3 Discussion
In general, BC-based methods struggle with generalization and modeling multi-modal action distributions. To overcome these limitations, diffusion models have recently attracted increasing attention. They can be employed either as decision models that directly generate action sequences [32] or as high-level strategy models that guide the action generation process [60]. In both settings, diffusion models have shown promising performance and improved flexibility over conventional BC methods.
III-B Inverse Reinforcement Learning
III-B1 Description
Inverse Reinforcement Learning (IRL) inverts the conventional RL framework, which focuses on learning a policy to maximize a predefined reward function. Instead, IRL aims to infer the underlying reward function that best explains a set of expert demonstrations.
Formally, IRL estimates a reward function $R(s,a)$ that best aligns with the demonstrated state-action pairs $\mathcal{D}=\{\tau_{1},\tau_{2},\dots,\tau_{N}\}$, where $\tau_{i}=\{(s_{0},a_{0}),(s_{1},a_{1}),\dots,(s_{t},a_{t})\}$. It is assumed that these demonstrations are generated by an expert following an optimal or near-optimal policy. The IRL problem is typically formulated within a finite Markov Decision Process, defined as $\mathcal{M}=\langle\mathcal{S},\mathcal{A},T,R,\gamma\rangle$, where $\mathcal{S}$ and $\mathcal{A}$ are the state and action spaces, $T(s^{\prime}|s,a)$ is the state transition probability, $R(s,a)$ is the reward function, and $\gamma\in[0,1]$ is the discount factor. IRL often represents the reward function as a linear combination of feature functions:
$R(s_{t},a_{t})=w^{\top}\phi(s_{t},a_{t})$ (2)
where $\phi(s,a)$ is a feature vector and $w$ is a learnable weight vector. The expected feature counts under a policy $\pi$ are defined as:
$\mu_{\phi}(\pi)=\sum_{t=0}^{\infty}\gamma^{t}\psi^{\pi}(s_{t})\phi(s_{t},a_{t})$ (3)
where $\psi^{\pi}(s)$ denotes the state-action visitation frequency:
$\psi^{\pi}(s)=\psi_{0}(s)+\gamma\sum_{s^{\prime}\in S}T(s^{\prime}|s,a)\psi^{\pi}(s^{\prime})$ (4)
IRL is particularly advantageous in dexterous manipulation scenarios, where manually defining a reward function is often challenging or impractical. IRL has demonstrated effectiveness in various dexterous manipulation tasks, including dexterous grasping, assembly, and manipulation in dynamic and uncertain environments.
III-B2 Research Progress
Recent studies have leveraged IRL frameworks to tackle complex dexterous manipulation tasks. Orbik et al. [63] first advance IRL for dexterous manipulation by introducing reward normalization, task-specific feature masking, and random sample generation. These techniques effectively mitigate reward bias toward demonstrated actions and enhance learning stability in high-dimensional state-action spaces, leading to better generalization across unseen scenarios. Building upon the need for efficient learning in such high-dimensional settings, Generative Causal Imitation Learning [64] improves the sample efficiency of IRL by integrating maximum entropy modeling with adaptive sampling strategies. By leveraging nonlinear function approximation through neural networks, the proposed method enables expressive cost function learning while handling unknown system dynamics. To further incorporate user feedback into the learning process, ErrP-IRL [65] integrates error-related potentials [66] with IRL. This approach assigns trajectory weights based on users’ cognitive responses, which are then used to iteratively refine a reward function represented as a sum of radial basis functions.
Beyond human feedback, recent works have explored learning reward functions from large-scale, unstructured demonstrations. GraphIRL [67] extracts task-specific embeddings from diverse video demonstrations. By modeling object interactions as graphs and performing temporal alignment, GraphIRL learns transferable reward functions without requiring explicit reward design or environment correspondence, enabling cross-domain manipulation capabilities. To further improve policy precision, Naranjo-Campos et al. [68] propose to integrate IRL with Proximal Policy Optimization [45]. Their method incorporates expert-trajectory-based features and a reverse discount factor to address feature vanishing issues near goal states, thereby improving the robustness of the learned policies. More recently, Visual IRL [69] extends the scope of IRL to human-robot collaboration tasks. It employs adversarial IRL to infer reward functions from human demonstration videos and introduces a neuro-symbolic mapping that translates human kinematics into robot joint configurations. This approach not only ensures accurate end-effector placement but also preserves human-like motion dynamics, facilitating natural and effective robot behavior in dexterous manipulation tasks.
III-B3 Discussion
In summary, IRL has demonstrated significant potential for dexterous manipulation tasks. By inferring the underlying reward function from expert demonstrations, IRL enables robots to generalize complex behaviors and adapt to diverse environments without the need for manually designed reward functions. This capability is particularly valuable in dexterous manipulation scenarios where reward specification is challenging or impractical [70, 71]. Despite these promising developments, state-of-the-art IRL methods still face several limitations. One of the primary challenges lies in accurately estimating reward functions, particularly in environments with high-dimensional action spaces or sparse feedback signals. Furthermore, IRL methods often rely on large amounts of expert demonstration data, which poses practical constraints due to the high cost and time required for data collection [70, 72].
III-C Generative Adversarial Imitation Learning
III-C1 Description
Generative Adversarial Imitation Learning (GAIL) extends the Generative Adversarial Network (GAN) framework [102] to the domain of imitation learning. It formulates the imitation process as a two-player adversarial game between a generator and a discriminator. The generator corresponds to a policy $\pi$ that aims to produce behavior that closely resembles expert demonstrations, while the discriminator $D(s,a)$ evaluates whether a state-action pair $(s,a)$ originates from the expert data $M$ or is generated by $\pi$.
Specifically, GAIL minimizes the Jensen-Shannon divergence between the state-action distributions of the expert and the generator. The discriminator is trained to maximize the following objective:
$\arg\min_{D}-\mathbb{E}_{d^{M}(s,a)}[\log D(s,a)]-\mathbb{E}_{d^{\pi}(s,a)}[\log(1-D(s,a))]$ (5)
where $d^{M}(s,a)$ and $d^{\pi}(s,a)$ denote the state-action distributions of the expert and the generator, respectively. The generator’s policy $\pi$ is optimized using RL with a reward signal derived from the discriminator:
$r_{t}=-\log(1-D(s_{t},a_{t}))$ (6)
Through this adversarial training process, GAIL effectively learns complex behaviors from expert demonstrations without explicitly recovering the reward function.
III-C2 Research Progress
GAIL has been widely adopted in dexterous manipulation. However, its effectiveness heavily relies on the quality and availability of expert demonstrations, which are often labor-intensive to collect and prone to inconsistencies [73, 74]. Such discrepancies arise from factors like collector biases [75], expert errors, noisy data, non-convex solution spaces, and suboptimal strategies [76]. Additionally, data scarcity further limits learning efficiency and policy robustness.
To address these challenges, various GAIL extensions have been proposed. HGAIL [77] employs hindsight experience replay to synthesize expert-like demonstrations without requiring real expert data. AIL-TAC [76] introduces a semi-supervised correction network to refine noisy demonstrations. GAIL has also been used in sim-to-real transfer [78], reducing the dependence on real-world expert data. Nevertheless, GAIL still suffers from mode collapse, where learned policies capture only a narrow range of behaviors, and gradient vanishing issues when the discriminator overpowers the generator. To mitigate these problems, RIDB [79] incorporates variational autoencoders to learn semantic policy embeddings and enable smooth interpolation across behaviors. WAIL [80] leverages the Wasserstein GAN framework [81] to improve training stability and reduce mode collapse. DIL-SOGM [82] further introduces a self-organizing generative model to capture multiple behavioral patterns without requiring encoders. In parallel, several works improve GAIL’s robustness under imperfect demonstrations. GA-GAIL [83] employs a second discriminator to identify goal states, enhancing policy learning from suboptimal data. RB-GAIL [84] integrates ranking mechanisms and multiple discriminators to model diverse behavior modes while leveraging generated experiences.
In addition to addressing the quality and availability of expert demonstrations, recent studies have sought to improve the performance of GAIL-based dexterous manipulation methods in other aspects. For instance, TRAIL [85] introduces constrained discriminator optimization to prevent the discriminator from focusing on spurious, task-irrelevant features such as visual distractors, thereby preserving meaningful reward signals and enhancing task performance. In the context of human imitation, Antotsiou et al. [86] combine inverse kinematics and particle swarm optimization with GAIL to mitigate sensor noise and domain discrepancies, enabling robots to autonomously grasp objects in simulation environments. Furthermore, P-GAIL [87] incorporates entropy-maximizing deep P-networks into GAIL to improve policy learning in deformable object manipulation tasks.
III-C3 Discussion
Although several extensions have addressed specific challenges of GAIL in dexterous manipulation, it still inherits the fundamental limitations of adversarial training. In particular, GAIL often suffers from training instability and faces difficulties in scaling to high-dimensional action spaces.
III-D Hierarchical Imitation learning
III-D1 Description
Hierarchical Imitation Learning (HIL) is an imitation learning framework designed to address complex tasks by decomposing them into a hierarchical structure. HIL typically adopts a two-level hierarchy, where the high-level policy is responsible for generating a sequence of sub-tasks or primitives based on the current state and task requirements, and the low-level policy executes sub-tasks to achieve the overall objective. This hierarchical decomposition enables handling long-horizon and complex tasks more effectively by separating decision-making and control.
Mathematically, the high-level policy $\pi_{h}$ selects a primitive $p_{i}$ from a predefined set of primitives $\{p_{1},p_{2},\dots,p_{K}\}$:
$\pi_{h}(s_{t})=p_{i}$
where $i\in\{1,2,\dots,K\}$. The corresponding low-level policy $\pi_{p_{i}}$ then generates the action to execute the selected primitive:
$a_{t}=\pi_{p_{i}}(s_{t})$
The overall objective of HIL is to minimize the cumulative loss function $\mathcal{L}(\pi)$, which explicitly reflects the hierarchical structure of the policy by jointly optimizing both the high-level decision-making and the low-level control execution to achieve effective task decomposition and coordination:
$\mathcal{L}(\pi)=\sum_{t=1}^{T}\mathbb{E}_{(s_{t},a_{t})\sim\pi}\left[\ell(s_{t},a_{t})\right]$ (7)
where $\ell(s_{t},a_{t})$ represents the immediate loss at time step $t$.
The parameters of the high-level and low-level policies in HIL are typically determined through three approaches: (1) Learning from demonstrations, which utilizes expert demonstrations to train both levels of policies; (2) Optimization, which applies RL or other optimization methods to refine the policies; and (3) Manual tuning, which manually adjusts policies during the initial stages or for specific task requirements. A key advantage of HIL is its ability to reduce the complexity of direct action-space search by decomposing tasks into hierarchical structures. This decomposition not only improves learning efficiency, particularly in long-horizon tasks, but also enhances generalization and task success rates by enabling optimization at multiple levels.
III-D2 Research Progress
In recent years, HIL has achieved significant progress in robotics, particularly in task decomposition and skill generalization. CompILE [88] enhances generalization in complex environments by decomposing tasks into independent sub-tasks. This compositional approach sets the foundation for subsequent research, particularly for tasks involving long temporal sequences. Building on this, Xie et al.[89] apply HIL to dual-arm manipulation and introduce the HDR-IL framework, which decomposes tasks into multiple motion primitives and employs graph neural networks to model object relationships.
HIL has also been successfully applied to various dexterous manipulation tasks. ARCH, proposed by Sun et al.[90], presents a framework tailored for long-horizon, contact-rich robotic assembly. It combines a low-level library of predefined skills with imitation learning for high-level policy, enabling efficient composition and adaptation of skills to handle complex, high-precision tasks. XSkill, introduced by Xu et al.[91], further extends HIL to cross-modal skill discovery. Their framework demonstrates how learning skill mappings across different modalities allow robots to generalize to diverse environments. Additionally, Wan et al. [92] propose LOTUS, which focuses on maintaining skill continuity and stability in dynamic environments by decomposing tasks into continuous sub-tasks and enabling policy adaptation across varying conditions. To efficiently train both high-level and low-level policies, recent works have explored the use of play data. Wang et al.[93] propose MimicPlay, which leverages unstructured human play interactions to guide robot manipulation. This approach learns high-level latent plans from play data and uses them to train a low-level visuomotor controller with only a small amount of demonstrations. Similarly, Lin et al.[94] introduce H2RIL, which utilizes cross-domain demonstrations by extracting interaction-aware skill embeddings from task-agnostic play data. These embeddings are aligned with human videos via temporal sequence contrastive learning, enabling the system to generalize to novel, composable tasks and adapt to out-of-distribution scenarios.
III-D3 Discussion
In summary, HIL has demonstrated significant advantages in task decomposition, skill generalization, and handling long-horizon tasks. By introducing hierarchical structures and multi-level control strategies, substantial progress has been made across various dexterous manipulation tasks. However, current research struggles with achieving adaptability in cross-modal skill generalization and ensuring model robustness and continuity in dynamic environments. In particular, when task environments change substantially, rapidly adapting and optimizing existing skill libraries to enable smooth task transitions remains an open problem.
III-E Continual Imitation Learning
III-E1 Description
Continual Imitation Learning (CIL) is an approach that integrates continual learning with imitation learning, aiming to enable agents to continuously acquire and adapt skills by imitating expert behaviors in dynamically changing environments. Specifically, in the initial phase, the agent learns fundamental skills from expert demonstrations. In subsequent phases, the agent incrementally accumulates knowledge, adapts to new tasks or environments, and mitigates the risk of forgetting previously acquired skills.
In CIL, the policy $\pi$ is optimized by minimizing the cumulative imitation loss across all previously encountered tasks. The objective function is defined as:
$\mathcal{L}(\pi)=-\sum_{i=1}^{t}\lambda(i)\,\mathbb{E}_{(s^{(i)},a^{(i)})\sim\rho^{(i)}_{\text{exp}}}\left[\log\pi\left(a^{(i)}\mid s^{(i)}\right)\right]$ (8)
where $\lambda(i)$ assigns a weight to each of the $t$ tasks, and $\rho^{(i)}_{\text{exp}}$ denotes the distribution of expert state-action pairs. The core objective of CIL is to continuously refine the policy $\pi$ using new demonstrations while preserving performance on previously learned tasks. This is particularly challenging, as it requires balancing the acquisition of new skills without compromising the proficiency of previously acquired ones.
III-E2 Research Progress
CIL has emerged as a key research direction in robotics, with the goal of enabling robots to continually acquire new skills from task demonstrations while mitigating catastrophic forgetting of previously learned tasks. To this end, various approaches have been especially proposed for dexterous manipulation.
Early studies in CIL focus on enabling robots to switch between tasks without compromising previously acquired skills, laying the groundwork for subsequent research [95]. However, these approaches often require substantial storage and computational resources, limiting their practicality in real-world applications. To address these limitations, researchers propose task-specific adapter structures, introducing lightweight adapters to facilitate seamless task switching [96]. This method effectively reduces storage overhead and improves adaptability to new tasks, though its performance tends to decline when tasks differ significantly. Further advancements explore the use of unsupervised skill discovery to enhance adaptability [92]. By dynamically generating new skills and integrating them into the robot’s existing skill set, this approach improves the robot’s ability to handle complex and evolving tasks. While promising results have been demonstrated in simulation, the effectiveness and generalization of these skills in real-world scenarios remain to be validated.
Further advancements in CIL have introduced the concept of learning unified policies through behavior distillation [97]. Unlike earlier works, this approach addresses the challenge of multi-task learning by employing a single policy structure, thereby eliminating the need for additional adapters when new tasks are introduced. However, designing such unified policies remains challenging, particularly in maintaining performance on previously learned tasks while accommodating new ones. In parallel, the emergence of GAN has inspired the development of CIL-based Deep Generative Replay (DGR)[98], which enables robots to continually acquire new skills by generating synthetic task trajectories. This approach alleviates the need to store past task data or recreate previous environments, effectively reducing storage overhead. Nevertheless, ensuring the quality and consistency of the generated trajectories remains an open challenge. Additionally, self-supervised learning has been explored to extract skill abstractions and expand the applicability of CIL to complex control tasks[99]. This method demonstrates that, even in the absence of explicit task demonstrations, robots can continually acquire new skills through self-supervision.
III-E3 Discussion
In summary, current research in CIL for dexterous manipulation mainly focuses on effective multi-task learning, the application of DGR techniques, and self-supervised skill abstraction. Through these methods, researchers have addressed some of the core challenges of continual learning, particularly in task switching and skill integration. However, significant challenges remain for practical deployment. The quality and consistency of generated data are not yet optimal, which may impact the accuracy of decision-making and task execution. Additionally, although some approaches have reduced storage and computational demands, resource consumption remains a limiting factor in more complex scenarios. Furthermore, the generalization capabilities of current methods, particularly in handling diverse tasks and dynamic environments, are still insufficient for real-world applications.
IV End Effectors for Dexterous Manipulation
An end effector is a component at the tip of a robotic manipulator that directly interacts with the environment to perform tasks. In dexterous manipulation, end effectors are typically categorized into two-fingered grippers, multi-fingered anthropomorphic hands, and three-fingered robotic claws. This section introduces these three types in order, highlighting their design principles, advantages, and trade-offs.
IV-A Two-fingered Traditional Gripper
Two-fingered grippers are widely used for their reliability, simplicity, and ease of control. Typically driven by a single actuator with one degree of freedom (DoF), they are cost-effective and suitable for repetitive tasks requiring consistency [103]. For example, [104] demonstrated a Franka robot with such a gripper performing tasks like setting a breakfast table. Similarly, Kim et al. [105] used a two-fingered gripper for behavior cloning with gaze prediction, and a tendon-driven variant in [106] showed the ability to grasp diverse household objects.
Recent works have further extended gripper capabilities through large-scale imitation datasets such as MIME [107], RH20T [108], Bridge Data [109], and Droid [110]. Dual-arm systems also enhance manipulation by coordinating two grippers. For instance, [111] achieved banana peeling through dual-action imitation learning. More complex tasks such as shrimp cooking, cloth folding, and dishwashing were demonstrated in Mobile ALOHA [112] and UMI [113].
Despite these advances, two-fingered grippers remain fundamentally limited in dexterous manipulation, which requires within-hand object reconfiguration [114]. Their simple structure and lack of internal DoFs restrict post-grasp adjustments [103]. Furthermore, morphological differences from the human hand hinder learning from demonstrations and prevent replication of human-like in-hand movements [115].
IV-B Multi-fingered Anthropomorphic Hand
To overcome the dexterity limitations of two-fingered grippers, robotic hands with human-like morphology have been widely developed. These anthropomorphic hands are better suited for interacting with objects and environments designed for humans [116]. They can be typically classified by transmission mechanisms—tendon-driven, linkage-driven, direct-drive, and hybrid systems—which fundamentally affect their performance characteristics [117, 118].
Figure 3: Examples of multi-fingered anthropomorphic hands: (a) Shadow Dexterous Hand[119]; (b) Awiwi Hand [120]; (c) Biomimetic Hand [121]; (d) ILDA Hand [117]; (e) Hu et al.’s robotic hand [118]; (f) INSPIRE-ROBOTS RH56 Dexterous Hand[122]; (g) OYMotion OHand[123]; (h) PUT-Hand [124]; (i) Allegro Hand[125]; (j) Faive Hand [126]; (k) Tesla Optimus Hand[127]; (l) Utah/M.I.T. Dexterous Hand [128].
IV-B1 Tendon-driven Approach
Tendon-driven hands use cable transmissions to actuate joints, mimicking human tendons. This design allows compact structure, multiple DoFs, and high dexterity, making it a common choice in anthropomorphic hand development. To accommodate high DoFs, actuators are often remotely located in the forearm.
Representative examples include the Utah/MIT Hand [129] (see Fig. 3(l)), the Shadow Dexterous Hand [130] (see Fig. 3(a)), and the Awiwi Hand [131, 132, 133, 134] (see Fig. 3(b)), all of which adopt antagonistic tendon routing for biomimetic motion. The FLLEX Hand [135, 136] and Faive Hand [137] (see Fig. 3(j)) with rolling contact joints demonstrate robustness and ball-rolling manipulation, respectively. Other typical tendon-driven hands include the Robonaut R2 Hand [138], Valkyrie Hand [139, 140], UB Hand [141, 142, 143], DEXMART Hand [144, 145], iCub Hand [146], and Biomimetic Hand [147] (see Fig. 3(c)).
While remote actuation reduces hand weight, it introduces friction and tendon wear due to long transmission paths. To address this, some designs embed all actuators within the palm. Examples include the DEXHand [148], SpaceHand [149], CEA Hand [150], and OLYMPIC Hand [151], which prioritize modularity and compact integration. Commercial designs like DexHand 021 [152], Tesla Optimus Hand [153] (see Fig. 3(k)) and PUDU DH11 Hand [154] also follow this approach.
Despite their advantages in dexterity and anthropomorphism, tendon-driven hands face challenges such as friction loss [155, 156], end termination [133, 157, 158], tendon creep and wear [159, 160, 161], which impact durability and reliability. As a result, most remain within research settings, with limited deployment in real-world industrial applications.
IV-B2 Linkage-driven Approach
Linkage-driven hands use rigid mechanical linkages to control joint motion, offering high precision, repeatability, and robustness. Compared to tendon-driven designs, they generally provide fewer DoFs but benefit from simpler, more reliable actuation. As a result, most commercial prosthetic and robotic hands adopt this mechanism. Due to space constraints and the demand for compactness, most linkage-driven fingers are actuated by a single motor and fall into two main categories: one-DoF coupled and multi-DoF underactuated types [162].
In the one-DoF type, joints are mechanically coupled, so preshaping remains fixed during flexion. Typical designs include the S-finger with inverse four-bar coupling [163] and the humanoid hand by Liu et al. using two four-bar linkages per finger [164]. Similar configurations appear in hands like the INSPIRE-ROBOTS RH56 [122] (see Fig. 3(f)), Bebionic Hand [165, 166], BrainRobotics Hand [167] and OYMotion OHand [123] (see Fig. 3(g)).
In contrast, underactuated fingers can adapt to contact forces, enhancing grasp adaptability. Examples include the Southampton Hand with a Whiffle tree mechanism [168], the LISA Hand using linkage-based self-adaptation [169]; MORA HAP-2 Hand [170], AR Hand III [171], and Cheng et al.’s prosthetic hand [172] with multi-bar or four-bar adaptive linkages.
While most designs use one motor per finger, a few incorporate multiple actuators for higher dexterity. The ILDA Hand [117] (see Fig. 3(d)) employs three motors per finger with combined PSS/PSU chains and four-bar linkages, achieving workspace and fingertip force comparable to human hands. Similar high-DoF linkage designs appear in the AIDIN ROBOTICS Hand [173] and the RY-H1 Hand [174].
IV-B3 Direct-driven Approach
Direct-drive hands eliminate intermediate transmission by connecting actuators directly to joints. This simplifies mechanical structure while still allowing for high actuatable DoFs, similar to tendon-driven designs.
Representative examples include the OCU-Hand [175] with 19 DoFs, where most joints are individually driven by embedded DC motors, and the TWENDY-ONE hand [176], which achieves 13 DoFs via joint-level motor placement. The KITECH-Hand [177], Allegro Hand [125] (see Fig. 3(i)) and LEAP Hand [178] adopt modular finger designs, integrating motors directly into the phalanges. The LEAP Hand also introduces a novel universal abduction-adduction motor configuration for enhanced MCP joint flexibility.
While direct drive offers high control precision and responsiveness, it introduces potential drawbacks such as increased mass, rotational inertia, and finger bulkiness, which may hinder agility in fine manipulation tasks. These limitations partly explain why most direct-drive hands adopt a four-finger configuration.
IV-B4 Hybrid-transmission Approach
In addition to the transmission types introduced above, many anthropomorphic hands adopt hybrid schemes to integrate the advantages of different approaches.
For example, the DLR/HIT Hand II [179] and NAIST Hand [180] use modular fingers with a combination of motors, belts, gears, tendon or linkage systems. The MCR-Hand series [181, 182] utilize a linkage-tendon mixed transmission system to achieve compactness and high functionality. Adab Mora Hand [183], LEAP Hand V2 (DLA Hand) [184] and Hu et al.’s hand [118] (see Fig. 3(e)) also integrate multiple transmission elements within fingers to enhance overall performance and adaptability.
Other hybrid designs explicitly differentiate mechanisms across fingers to match specific functional needs. The PUT-Hand [124] (see Fig. 3(h)), for instance, combines a direct-drive thumb, linkage-driven index/middle fingers, and tendon-driven ring/little fingers. Similarly, the MPL v2.0 Hand [185], Tact Hand [186], and Six-DoF Open Source Hand [187] adopt direct or geared actuation for the thumb while using tendons, linkages or timing belts for other fingers. Hands developed by Owen et al. [188], Ryu et al. [189], and Ke et al. [190] follow a similar approach, implementing thumb-specific hybrid strategies to enhance opposability and dexterity.
IV-C Three-fingered Robotic Claw: A Trade-off Solution
The diversity of anthropomorphic hand designs largely stems from a fundamental trade-off between mechanical simplicity and dexterous capability [191]. While the human hand has over 20 DoFs [192, 193, 194, 195, 196, 197, 198], replicating this complexity mechanically remains impractical. Higher dexterity often increases structural and control complexity, cost, and susceptibility to failure [199, 200, 201], limiting the feasibility of high-DoF hands in real-world applications [202, 203].
To mitigate these issues, several simplification strategies are adopted: underactuation with elastic components [200, 201, 130, 127, 168, 169], reducing non-essential DoFs or phalanges [127, 122, 165, 123], or even omitting a finger entirely [129, 139, 148, 176, 178, 125]. These approaches highlight the challenge of maximizing functionality within practical constraints.
Figure 4: Examples of three-fingered robotic claws: (a) DEX-EE [204]; (b) BarrettHand [205]; (c) i-HY Hand [199]; (d) DoraHand [206].
As a compromise between the minimalistic two-fingered gripper and complex multi-fingered anthropomorphic hands, the three-fingered robotic claw offers a functional middle ground. Though not anatomically human-like, three fingers are sufficient for executing common grasp types such as cylindrical and spherical power grasps [199], and can support a subset of in-hand manipulation tasks.
Numerous three-fingered claws have demonstrated impressive capabilities. For example, Shadow’s DEX-EE [204] (see Fig. 4(a)) and the TRX Hand [207] exhibit high robustness and dexterity. The BarrettHand [208] (see Fig. 4(b)) achieves adaptive grasping through underactuation. Tendon-driven designs like the i-HY Hand [199] (see Fig. 4(c)) and Model O [209] enable actions such as pivoting and precision transitions. Systems such as DClaw [210] and TriFinger [211] are capable of performing fine manipulation tasks via reinforcement learning. Other novel architectures include linkage-based [212], motor-multiplexed [213], and link-belt-integrated claws [214], each offering different characteristics.
Three-fingered claws such as the DoraHand [203] (see Fig. 4(d)), SARAH [215], D’Manus [216], and Kinova Jaco’s claw [217] further demonstrate the practicality and versatility of this design choice in both research and assistive applications.
V Teleoperation Systems and Data Collection
Teleoperation systems provide a robust interface for human-robot collaboration, benefiting from directly making robot behaviors comply with human-level intelligence, which refers to human-in-the-loop. This approach is highly intuitive since humans’ extensive knowledge and experience empower them to make informed judgments on diverse tasks across complex scenes and to promptly adjust strategies in response to feedback. Due to this usability, teleoperation is widely applied in various fields. Additionally, by collecting data on the robot’s states and corresponding actions during teleoperation, datasets can be constructed to perform end-to-end imitation learning.
V-A Teleoperation Systems for Dexterous Manipulation
A typical teleoperation system consists of two main components: the local site and the remote site, as demonstrated in Fig. 5. The local site includes a human operator and a suite of interactive I/O(input/output) devices. The output devices provide real-time status about the remote robot and its surrounding environment, while the input devices allow the operator to issue commands in diverse forms, thereby controlling the remote robot’s actions. The remote site primarily contains the robot itself, which is equipped with various sensors to gather perceptions of its state and the surrounding environment. Upon receiving teleoperation commands from humans, the robots can perform the corresponding actions and complete tasks.
Figure 5: Teleoperation Frameworks and Commonly Used Devices. (a) mocap gloves (b) VR controllers (c) joystick (d) camera (e) exoskeleton (f) dexterous hand (g) dual-arm robot (h) robot arm.
To accurately convey human operators’ intentions to robotic systems, previous works have employed a wide range of human-robot interaction devices. Human operators with work experience can easily identify the current state of a robot through the image, however, accurately translate human instructions to robot actions remains a challenge. Some traditional controllers act on this: 1) joysticks [218] 2) haptic devices [219]; However, manipulation tasks often involve delicate movements and complex interactions, such as grasping, moving, and positioning small or irregularly shaped objects. These tasks necessitate devices that can offer dexterous interfaces to ensure the safety and efficacy of the robot’s actions. Precision and real-time feedback are crucial. Commonly used devices include: 1) cameras [18, 220, 17, 19, 221, 222, 223, 54, 20]; 2) mocap gloves [224, 225, 226, 227, 228, 229]; 3) VR/AR controllers [230, 231, 14, 232, 233, 54, 234, 27, 235, 236, 237, 238]; 4) exoskeletons and bilateral systems [239, 240, 241, 53, 242, 243].
V-A1 Vision-based Teleoperation Systems
Recently, advancements in computer vision have led to the development of vision-based teleoperation systems. However, their accuracy in capturing hand movements is often compromised by factors such as occlusion, lighting, resolution, background, and inaccurate 3D estimation issues. Several methods have been proposed for robust hand pose estimation and reliable mapping to the robot end-effector. Li et al. [222] developed a vision-based teleoperation system by training TeachNet on pairs of images of human hands and simulated robots to form mappings between a human hand and a robotic Shadow Hand in the latent space. Dexpilot [18] utilized a calibrated multi-camera system to estimate hand poses to teleoperate an Allegro Hand. Riemannian Motion Policies (RMPs) are employed to compute the Cartesian pose of the hand, facilitating hand-arm motion control. Subsequent approaches, such as Robotic Telekinesis [17] and DIME [54], which simplified requirements to a single RGB camera, thus reducing the need for calibration. This is achieved through a general mapping method between humans and robots that have different kinematic structures. Additionally, Robotic telekinesis [17] adjusts the position and orientation of the end-effector relative to its base using the relative position and direction of the human wrist to the torso, enabling the teleoperate both arm and hand. However, these methods still suffer from occlusion issues due to the single fixed camera setting. To solve this problem, Transteleop [223] introduced a system that utilizes real-time active vision with a depth camera mounted on the end-effector of the remote UR5 robot arm. During teleoperation, this robot arm can reposition the camera to enhance its field of view and improve hand pose estimation accuracy.
The morphology discrepancy between the human hand and the robot hand might impede the operator from intuitively controlling the robot. To address this, Qin et at. [20] developed a user-friendly interface by constructing a customized robotic hand modeled after the specific shape of a human hand. Demonstrations performed with this customized robot hand can be directly transferred to any dexterous robot hand. AnyTeleop [19] proposed a solution to the self-occlusion problem by integrating images from multiple cameras, each offering different perspectives. To further enhance precision in observation, ACE [221] mounted the camera under the end-effector of the exoskeleton to maintain a clear view of the hands and wrists. MimicPlay [93] employs two calibrated cameras in different viewpoints to reconstruct 3D hand locations. The teleoperation data for the robot is collected using the RoboTurk system [244], which operates via an IMU-equipped smartphone.
V-A2 Mocap Gloves
Motion capture systems typically utilize stable hardware devices such as multi-camera setups with markers, IMU sensors, and RGB-D cameras. These devices are robust against changes in lighting, occlusion, and complex backgrounds. Mocap gloves collect human hand motion data directly via sensors, enabling ideal real-time performance and significantly improving data collection efficiency in teleoperation. Although motion capture gloves are expensive, they provide precise hand tracking [226]. Wang et al. [16] introduced DexCap, a portable motion capture system. It includes a mocap glove for accurate finger joint tracking, a single-view camera for 6-DoF wrist pose tracking, and an RGB-D LiDAR camera for observing the surrounding 3D environment. With this precise 3D hand motion data, the proposed DexIL system can effectively learn bimanual dexterous manipulation skills. The remote system features two Franka Emika robotic arms, each outfitted with a LEAP dexterous robotic hand. Similarly, Mosbach et al. [245] used the SenseGlove DK1, a force-feedback glove, to capture hand joint movements, with hand tracking facilitated by a camera mounted on a headset.
V-A3 VR/AR Controllers
VR devices typically include a head-mounted display, a tracking system, and input devices. The head-mounted display provides an immersive visual experience with high-resolution screens and head motion tracking. The tracking system captures the user’s movements to ensure that interactions in the virtual environment correspond to real-world actions. Input devices, such as controllers or gloves, facilitate user interaction within the virtual space. Zhang et al. [14] developed a teleoperation system using consumer-grade VR devices to control a PR2 robot. Following this, methods utilizing low-cost equipment [234, 246] demonstrated high-quality teleoperation through mixed reality. To streamline scene construction, Mosbach et al. [245] explored VR teleoperation in simulated environments for manipulation tasks. Recently, Bunny-VisionPro [235] equipped the Apple VisionPro with a haptic module to provide tactile feedback. Similarly, Open-television [247] used an active camera mounted on a humanoid robot to capture first-person stereo videos. This approach enhances the robot’s ability to perform precise and context-aware actions by providing a dynamic, real-time visual perspective similar to human vision. Lin et al. [248] introduced a low-cost teleoperation system, HATO, combining two Psionic Ability Hands for prosthetic use with UR5e robot arms. The system utilizes two Meta Quest 2 VR controllers with IMU sensors to capture hand spatial positions and orientations, translating controller inputs into multi-fingered hand poses.
V-A4 Exoskeleton and Bilateral Systems
The majority of aforementioned methods focus on manipulating the robot’s end-effector in task space in Cartesian coordinates. While setting the robot’s end-effector position is convenient, it has drawbacks. For robots with multiple degrees-of-freedom (multi-DoF), computationally demanding inverse kinematics (IK) calculations are required, which can be problematic in real-time control scenarios. These complexities may cause response delays and compromise operational precision. Furthermore, singularities in the motion trajectory may lead to indeterminate or nonexistent IK solutions, resulting in control failures. In the following sections, we examine studies that aim to synchronize human and robot movements in the joint space.
Exoskeletons are wearable devices that gather and analyze user motion data. Fabian [239] developed a lightweight exoskeleton, DE VITO for measuring human arm movements to teleoperate the mobile robot DE NIRO[249]. Meanwhile, AirExo [240] presents a framework for whole-arm dexterous manipulation adaptable to different robot arms using interchangeable 3D-printed components for robots divergent in morphology.
Another approach involves a bilateral framework, where movements of the leader robot are mirrored by the follower robot. Any resistance or force encountered by the follower is communicated back to the leader, enabling precision and tactile sensation tasks. Kim et al. [241] develops a controller with Denavit-Hartenberg (DH) parameters matching the teleoperated dual-arm robot, alongside a calibration method to reduce gravitational errors. For demonstrations without real robots, the controller uses force/torque (F/T) sensors identical to those in real robots to provide force feedback. Recently, ALOHA [53] utilizes structurally analogous robotic arms with identical joint spaces for teleoperation, employing cost-effective ViperX [250] aarms as follower robots and WidowX [251] of comparable size serve as leaders to enhance control capabilities
Expanding upon this concept, Mobile ALOHA [242] integrates the system with an Automated Guided Vehicle (AGV) to establish a whole-body teleoperation system. GELLO [243] has further reduced costs by replacing the real robotic arms at the local site with scaled kinematically equivalent 3D-printed parts and off-the-shelf motors, achieving one-to-one joint mapping.
Differing from specific robot methods, AnyTeleop [19] introduced a unified system supporting multiple robot arms and dexterous hands through a general motion retargeting method, which maps robot hands to human fingers. This system supports different arms by generating trajectories based on the estimated Cartesian end-effector pose. ACE [221] developed a cross-platform visual-exoskeleton teleoperation system compatible with diverse robot hardware, including various end-effectors such as grippers and multi-finger hands, offering flexibility. Its exoskeleton arm features high-resolution encoders for precise joint position readings, ensuring accurate end-effector tracking.
V-B Datasets and Benchmarks
V-B1 Datasets
MIME [252]is a large-scale dataset contains 8,260 human-robot demonstrations across 20 diverse tasks ranging from simple tasks like pouring to difficult tasks like stacking objects. It includes both videos of human demonstrations and kinesthetic trajectories for robots.
RH20T [253] dataset encompasses over 110,000 multimodal robotic manipulation sequences collected using intuitive teleoperation interfaces equipped with force-torque sensors and haptic feedback. It captures visual, tactile, audio, and proprioceptive data alongside human demonstration videos, facilitating one-shot imitation learning across diverse tasks and robotic configurations. This dataset aims to advance robotic skill acquisition in unstructured environments through enhanced task and motion planning.
BridageData [254] encompasses 7,200 demonstrations across 71 tasks in 10 environments. Leveraging diverse and complex manipulations primarily in kitchen settings. It aims to support broad skill generalization via cross-domain datasets. Following this, BridgeData V2 [255] expands the range of tasks and environments to foster more robust generalization and transfer capabilities in robots. With 60,096 robotic manipulation trajectories across 24 environments, it supports scalable robot learning with tasks varying from pick-and-place to complex manipulation, facilitating generalization across tasks, objects, and settings for multi-task and language-conditioned learning methods.
DRIOD [256] surpasses other datasets collected purely via human teleoperation with its unparalleled scene diversity and task variety. It is a diverse robot imitation learning dataset with 76,000 demonstration trajectories across 86 tasks and 564 scenes.
To address the problem that collecting extensive human demonstrations for imitation learning is always time-consuming and laborious. Several datasets apply data augmentation to demonstrations collected via human teleoperation. To illustrate, data augmentation is a technique that has been widely applied in various fields, particularly in computer vision and robotics. It is used to artificially expand the size and diversity of a training dataset by applying a series of transformations or modifications to the original data, which can improve the robustness and generalization capabilities of machine learning models trained on the augmented dataset.
RoboAgent [257] contains 7,500 trajectories collected through human teleoperation. It can be scaled up to approximately 98,000 trajectories to diversify data through semantic augmentations without extra human/robot cost. Similarly, CyberDemo [258] follows this approach. Researchers collect human demonstrations using teleoperation in both simulated and real-world environments and then implement extensive data augmentation to the demonstrations collected. By incorporating visual and physical variations in simulations, robots trained in them could obtain enhanced policy robustness and generalization ability.
Several other datasets leverage the demonstration generation system to enhance imitation learning by expanding the dataset from a limited number of human demonstrations. In MimicGen [259], over 50,000 demonstrations are created across 18 tasks from approximately 200 human demonstrations, it synthesizes full demonstrations of diverse scene configurations, object instances, and robot arms by adapting object-centric manipulation behaviors to new contexts through trajectory transformations based on known object poses, enabling the training for complex, long-horizon, and high-precision tasks.
Similarly, IntervenGen[260], introduced in 2024, also autonomously produces large sets of corrective interventions from minimal human input, enhancing policy robustness against distribution shifts. It specializes in generating interventional data to address policy mistakes, further reducing human effort and improving robustness.
DiffGen [261] basically follows the very same idea but integrates differentiable physics simulation, rendering, and vision-language models to generate realistic robot demonstrations from text-based instructions.
There are also datasets that focus on dexterous bimanual manipulations and hand-object interactions. For example, ARCTIC [262] contains 2.1 million videos with accurate 3D hand-object meshes and dynamic contact data, enabling the study of dexterous bimanual manipulation of articulated objects. It introduces novel tasks for consistent motion reconstruction and interaction field estimation, facilitating advanced research in hand-object interaction.
Additionally, DexGraspNet [263] featuring 1.32 million grasps of 5,355 objects by ShadowHand fills the blank of large-scale, diverse, and high-quality datasets for dexterous grasping. It includes over 200 diverse grasps per object, validated for physical stability in simulation, enabling more effective imitation learning and benchmarking of robotic manipulation algorithms to achieve human-like dexterity and grasping capabilities.
The OAKINK2 dataset [264] comprises 627 sequences of bimanual object manipulation, featuring 4.01 million frames from multi-view captures and detailed pose annotations for human bodies, hands, and objects.
VI Challenges and Future Directions in Imitation Learning-Based Dexterous Manipulation
Imitation learning-based dexterous manipulation poses unique challenges due to the inherent complexities of both imitation learning and dexterous control. Despite significant advancements over the past decade, several challenges hinder its human-level dexterity and real-world applicability. This section discusses these challenges from multiple perspectives and explores future research directions.
VI-A Data Collection and Generation
Data collection and generation for imitation learning-based dexterous manipulation poses several challenges, including heterogeneous data fusion, data diversity, high-dimensional data sparsity, and data collection costs:
VI-A1 Heterogeneous Data Fusion
Dexterous manipulation relies on multi-modal sensory inputs (e.g., visual, tactile, proprioceptive, and force), each with varying sampling rates, noise characteristics, and spatial-temporal resolutions, making data integration and synchronization challenging. Moreover, differences in embodiments and gripper designs introduce additional complexities. For example, demonstrations collected with one robotic hand may not directly generalize well to another due to variations in kinematics, actuation mechanisms, and sensor placements. Addressing these challenges requires (1) multi-modal alignment techniques to improve sensor fusion and (2) cross-embodiment learning frameworks for better transferability across robotic platforms and embodiments.
VI-A2 Data Quantity, Quality, and Diversity
Ensuring sufficient data quantity, quality, and diversity is challenging because collecting expert demonstrations for dexterous tasks at scale is labor-intensive and expensive. Even small variations in object properties, task conditions, or environmental factors can significantly affect manipulation policies, making it difficult for imitation learning models to generalize. Future research should explore synthetic data augmentation, domain randomization, and generative models to efficiently generate diverse training datasets. Scalable and automated data collection methods, such as crowdsourced teleoperation, where multiple users remotely control robots to provide varied demonstrations, and self-supervised learning, where robots autonomously collect and label data through interaction and feedback, can further mitigate data collection bottlenecks. Additionally, establishing standardized data collection protocols and defining robust evaluation metrics for data quality and diversity will be essential for ensuring consistency and reliability.
VI-A3 High-Dimensional Data Sparsity
Data sparsity in high-dimensional action spaces limits the effectiveness of learned policies, as dexterous manipulation requires precise finger coordination, force regulation, and contact-rich interactions that demonstrations alone struggle to capture comprehensively. Hierarchical representation learning can potentially mitigate this challenge by structuring high-dimensional action spaces into more learnable subspaces. In dexterous manipulation, decomposing control policies into hierarchical levels—such as low-level motor commands, mid-level grasp strategies, and high-level task affordances—allows models to extract structured representations, improving learning efficiency and reducing dependence on large-scale demonstrations.
Reinforcement learning fine-tuning further complements imitation learning by refining dexterous manipulation policies beyond demonstrated behaviors. Fine-tuning in simulation enables robots to explore variations in object properties, task conditions, and environmental dynamics that may not be covered in the demonstration data. However, effective sim-to-real transfer techniques and high-fidelity physics engines are crucial to bridging the gap between simulated training and real-world execution.
VI-A4 Data Collection Costs
The high cost and complexity of data collection pose barriers to scaling imitation learning for dexterous manipulation. Traditional methods often require specialized motion capture systems, high-precision force sensors, and complex teleoperation setups, which are expensive, labor-intensive, and impractical for large-scale data acquisition. Reducing these barriers requires the development of low-cost, scalable data collection methods, such as wearable sensor systems for capturing human demonstrations and shared autonomy techniques to minimize operator effort. Additionally, establishing standardized data collection protocols and collaborative data-sharing platforms can improve data accessibility and consistency across datasets.
While simulation provides a scalable solution for generating synthetic data in dexterous manipulation, several challenges limit its real-world effectiveness. First, achieving real-world fidelity remains difficult, as physics engines struggle to model contact dynamics, deformable objects, and high-resolution tactile feedback, leading to discrepancies between simulation and reality. Second, ensuring sufficient data diversity is another challenge, as models trained in static or overly idealized environments often fail to generalize to unstructured real-world conditions, and while domain randomization can enhance robustness, excessive variation may reduce learning efficiency or introduce unrealistic artifacts. Third, the sim-to-real gap further complicates deployment, as policies trained in simulation often fail in real-world settings due to sensor noise, unexpected disturbances, and actuation discrepancies. While techniques such as domain adaptation, sim-to-real fine-tuning, and physics-based calibration can help mitigate these challenges, they require substantial computational resources and real-world validation, increasing deployment complexity.
VI-B Benchmarking and Reproducibility
The dependence on real-world hardware experiments and the variability in simulation environments pose significant challenges for benchmarking and reproducibility in imitation learning-based dexterous manipulation. Unlike computer vision or natural language processing, where large-scale datasets enable standardized evaluations, dexterous manipulation involves physical interactions, making consistent replication across research efforts difficult. Hardware dependency is a major obstacle, as reproducing results requires access to the same robotic platform, gripper design, sensor setup, and control software, which is often impractical in real-world experiments due to cost, availability, and proprietary constraints.
Simulation-based benchmarks offer a scalable alternative, but the lack of standardized simulation settings, computing environments, and evaluation protocols in physics-based simulators limits fair comparisons across studies. Variability in physics engine configurations, actuator models, contact dynamics, and material properties further exacerbates inconsistencies, making it difficult to establish reliable performance benchmarks and universally comparable evaluation metrics in dexterous manipulation research. Some studies rely on non-physics-based or simplified simulators, which focus on high-level task planning but neglect low-level contact physics modeling. While these environments provide visual realism and scalable training, they introduce a significant sim-to-real gap, failing to capture key aspects of dexterous manipulation, such as precise force interactions and object deformations.
Addressing these challenges requires standardized benchmarking frameworks and open-source datasets for both simulation and real-world experiments. In simulation, standardization should focus on consistent physics parameterization (e.g., contact dynamics, actuator models, material properties) and common environment representations to minimize discrepancies across different physics engines. For real-world experiments, benchmarks should incorporate multi-modal sensory recordings (e.g., RGB-D, tactile, proprioceptive data) and diverse task demonstrations across various robotic embodiments to ensure broader comparability. Additionally, establishing standard evaluation protocols across hardware platforms and physics-based simulators would enable more reliable performance comparisons across studies.
VI-C Generalization to Novel Setups
Generalizing imitation learning-based dexterous manipulation policies is challenging due to task and environment variability, adaptive learning limitations, sim-to-real transfer issues, and cross-embodiment adaptability:
VI-C1 Task and Environment Variability
Learning-based policies often struggle to extend beyond the specific demonstrations to new conditions. Variations in object shapes, sizes, weights, textures, and dynamic interactions, as well as unforeseen obstacles and workspace changes, can significantly degrade performance. Also, these policies may fail when faced with unseen task configurations that require adaptive behavior beyond the demonstrated distribution.
VI-C2 Adaptive and Continual Learning Frameworks
Traditional imitation learning models do not adapt to new tasks or environmental changes after training. This limitation leads to rigid behaviors that fail to improve with experience. Continual learning frameworks allow robots to learn incrementally from new data without catastrophic forgetting, while adaptive learning methods such as meta-learning and reinforcement learning fine-tuning enable policies to generalize to new conditions by leveraging prior experience. Additionally, uncertainty-aware models can dynamically adjust decision-making strategies based on real-time feedback, improving generalization in unstructured settings.
VI-C3 Sim-to-Real Transfer
While simulation provides a scalable and controlled environment for training dexterous manipulation policies, transferring these learned behaviors to real-world settings is challenging. Differences in contact dynamics, sensor noise, actuation delays, and material properties create a sim-to-real gap, causing trained models to perform inconsistently when deployed on real robots. Potential future research directions are improving the realism of physics simulations through differentiable physics engines, adaptive parameter tuning, and self-supervised real-to-sim refinement to better approximate real-world interactions. Additionally, leveraging hybrid learning approaches, where policies are pre-trained in simulation and fine-tuned with real-world corrections, can enhance transferability. Uncertainty estimation techniques can also be integrated to help models recognize and adapt to distribution shifts when deployed in unstructured real-world environments.
VI-C4 Cross-Embodiment Adaptability
Variability in robot embodiments, gripper designs, sensor configurations, and actuation dynamics pose significant challenges for generalization. A policy trained on one robotic hand may struggle to transfer to another due to differences in degrees of freedom, joint limits, contact dynamics, and control strategies. Even within the same robotic platform, inconsistencies arise from sensor noise, latency, and mechanical tolerances. To address this, morphology-agnostic policy learning can be explored, where models are trained across diverse robotic embodiments to develop transferable representations. Graph-based and latent-space embeddings of robot kinematics could help policies reason about different embodiments more effectively. Additionally, modular policy architectures, where separate components (e.g., perception, control, and adaptation modules) are fine-tuned independently, may enhance transferability. Another promising direction is meta-learning and few-shot adaptation, enabling robots to quickly adjust to new embodiments with minimal data, reducing the need for extensive retraining.
VI-D Real-Time Control
Dexterous manipulation presents significant computational challenges due to its high-dimensional action spaces and complex dynamics. Achieving real-time execution demands a delicate balance between accuracy and efficiency in terms of both software and hardware.
Efficient real-time control relies on algorithms capable of handling nonlinearities, contact dynamics, and feedback loops while maintaining stability and responsiveness. Model-based approaches, such as optimal control and Model Predictive Control (MPC), leverage system dynamics to generate control policies but often struggle with the complexities of dexterous manipulation. MPC, in particular, provides real-time adaptability through continuous optimization but imposes high computational demands, often requiring specialized hardware acceleration or dedicated edge computing to meet real-time constraints. In contrast, model-free reinforcement learning learns policies directly from data, bypassing the need for explicit system modeling. While reinforcement learning offers greater adaptability in high-dimensional, unstructured environments, it remains sample inefficient, prone to slow convergence, and challenging to stabilize, especially for real-time execution. A potential solution is designing hybrid control strategies that combine model-based control for stability with model-free learning for adaptability, improving efficiency without sacrificing robustness. Meanwhile, accelerated learning techniques, such as parallelized reinforcement learning training and meta-learning, could address sample inefficiency, enabling faster policy convergence.
Hardware architecture is also a key enabler of real-time dexterous manipulation, balancing computational power, latency, and energy efficiency. High-performance computing hardware (e.g., GPUs, TPUs, and FPGAs) is essential for complex model-based and learning-based control strategies but is often constrained by high power consumption and deployment costs. Edge computing and custom ASICs offer low-latency processing but may lack the computational capacity required for large-scale dexterous manipulation policy inference. Cloud computing facilitates large-scale training and high-fidelity simulations; however, real-time reliance on remote processing is limited by communication delays and network instability. Recent advancements in low-power AI accelerators, neuromorphic computing, and distributed edge-cloud architectures have the potential to enhance real-time processing while reducing latency and energy constraints.
VI-E Safety, Robustness, and Social Compliance
Ensuring safety, robustness, and social compliance is crucial for real-world dexterous robotics, requiring risk prevention, adaptive error recovery, and human-aware behavior for seamless integration.
Real-world dexterous manipulation presents significant challenges in error detection, recovery, and adaptability, requiring robots to operate reliably in dynamic and unstructured environments. Failure detection is complex due to sensor noise, occlusions, and unpredictable interactions, making it difficult to distinguish minor execution deviations from critical failures such as grasp failures or unexpected object motion. Once an error is detected, adaptive recovery strategies such as re-grasping and trajectory replanning must be computed in real-time while maintaining stability and task continuity. Future research should address two key aspects. First, large-scale failure datasets and standardized benchmarks are essential for improving data-driven recovery policies. The lack of diverse, labeled failure cases across various objects, tasks, and environments limits model generalization. Establishing comprehensive datasets and evaluation protocols for failure detection, uncertainty estimation, and recovery effectiveness would provide a foundation for training and benchmarking robust policies. Second, self-supervised learning for multi-modal anomaly detection could enable robots to autonomously refine their error detection capabilities. By leveraging visual, tactile, and proprioceptive feedback, robots could learn to recognize and anticipate failures in real-time, improving adaptability and robustness in dynamic environments.
Safety is equally critical for both the robot and its surrounding environment, including human users, particularly in real-world deployments where unpredictable interactions and dynamic conditions pose significant risks. In dexterous manipulation, safety considerations involve collision avoidance, force regulation, and compliance control, particularly when interacting with fragile objects or operating near humans. However, achieving these safety measures requires handling varying contact conditions, but sensor noise, occlusions, and data processing delays can reduce reliability. Additionally, while compliant actuators and soft robotic designs help mitigate impact forces, integrating these hardware safety mechanisms involves trade-offs between control precision, responsiveness, and durability.
Beyond technical safety, social compliance is crucial for real-world deployment while less explored, particularly in human-robot interaction settings. Robots must adhere to social norms, ethical guidelines, and human expectations to be perceived as trustworthy and acceptable. This includes adapting manipulation strategies to align with human workspaces, ensuring transparent behavior and predictability, and minimizing actions that could cause discomfort or disruption. However, existing manipulation frameworks lack awareness of social constraints and human preferences. To address this challenge, interactive learning paradigms, where robots refine their socially compliant manipulation strategies by learning from human corrections and preferences, offer a promising research direction. Complementing this, multi-modal human-robot interaction datasets that integrate verbal, visual, and non-verbal cues would enhance contextual understanding, enabling robots to better anticipate and respond to human needs. Furthermore, ensuring consistency and reliability in socially aware dexterous manipulation requires standardized benchmarks for social compliance, providing objective evaluation criteria to assess how well robots integrate into human-centered environments.
VII Conclusion
Imitation learning has shown significant promise in enabling robots to perform dexterous manipulation tasks with human-like skill and precision. By learning from human demonstrations, robots can acquire complex manipulation capabilities that are difficult to achieve through traditional programming methods. This survey has provided an overview of the current state-of-the-art in imitation learning-based dexterous manipulation, highlighting key techniques, applications, and challenges.
Despite the progress has been made, several challenges remain that hinder the practical deployment of these systems. Addressing issues related to data collection, generalization, real-time control, safety, and sim-to-real transfer is essential for advancing the field. Future research should focus on developing optimized imitation learning algorithms, enhancing human-robot collaboration, and integrating advanced sensory systems.
The future of dexterous manipulation holds great potential, with applications ranging from industrial automation to healthcare and service robotics. By continuing to push the boundaries of imitation learning and robotic manipulation, researchers and practitioners can pave the way for more capable, adaptable, and intelligent robotic systems. These advancements will not only improve the efficiency and safety of robotic tasks but also open up new possibilities for human-robot collaboration and interaction.
Acknowledgment
We would like to thank Qianyi Wang, Mingwu Liu, and Shouzheng Wang for their contributions to the completion of this survey.